OpenStudy (anonymous):

The intensity of an electromagnetic wave is 8 x 10^7 W/m^2. What is the amplitude of the magnetic field of this wave?

OpenStudy (chillout):

An electromagnetic wave is composed by E and H, where E is the electric field and H is the magnetic field. The amplitude is the peak value.

OpenStudy (chillout):

Which is found by using this equation: I = (c*Bmax²)/(2μ0), where c is the speed of light, μ0 is the permeability, I is the intensity of the electromagnetic wave and Bmax is the peak value. Just plug in the values and solve for Bmax!

OpenStudy (chillout):

Similarly, we can found the max value of the electric field by doing c x Bmax (or Hmax), where x denotes the cross product. Since they are perpendicular to each other (the angle is 90°), sin(x) is just 1.

OpenStudy (chillout):

find*

OpenStudy (anonymous):

wow! and for permeability is a constant or i have to use another formula?

OpenStudy (chillout):

It is a constant. µ0 = 4π x 10^-7 T•m/A

OpenStudy (anonymous):

mmmm... so is 8x10^7= (299792458)(Bmax)^2/2(4pi)10^-7

OpenStudy (anonymous):

mmmm... so is 8x10^7= (299792458)(Bmax)^2/2(4pi)10^-7

OpenStudy (anonymous):

intensity = E^2/uoC E^2 = 8*10^7 *2* 4pi*10^-7 * 3*10^8 E = 24.55*10^4 N/C

OpenStudy (anonymous):

Here! intensity = E^2/uoC E^2 = 8*10^7 * 4pi*10^-7 * 3*10^8 E = 17.36*10^4 N/C now since E/B = c B = E/c = 17.36*10^4/3*10^8 = 5.78 *10^-4 T is this right?

OpenStudy (kropot72):

\[I=8\times 10^{7}=\frac{1}{2 \mu _{0}}E _{m}B _{m}...........(1)\] \[E _{m}=cB _{m}..........(2)\] Substituting for Em in equation (1) gives \[8\times 10^{7}=\frac{1}{2 \mu _{o}}cB _{m}...........(3)\] Rearranging equation (3) we get \[B _{m}=\sqrt{\frac{8\times 10^{7}\times 2\times 4 \pi \times 10^{-7}}{3\times 10^{8}}}\ T\]

OpenStudy (anonymous):

so is 8.2 x 10^-4 ???

OpenStudy (kropot72):

Sorry, my bad. Equation (3) should be written \[8\times 10^{7}=\frac{1}{2 \mu _{0}}cB _{m}^{2}\] My calculation resulted in 8.185 * 10^-4 T for the amplitude of the magnetic wave.

OpenStudy (anonymous):

nice! :) thanks!

OpenStudy (kropot72):

You're welcome :)